OpenStudy (anonymous):

A jar has 6 blue balls and 4 yellow balls. If two are drawn without replacement, whats the probability that neither ball is yellow?

Parth (parthkohli):

That means that both balls or blue. \(\color{Black}{\Rightarrow \large {6 \over 10} \times {5 \over 9} }\)

OpenStudy (anonymous):

why wouldn't you use something like combinations?

Parth (parthkohli):

Why would you try to use combinations here? :)

OpenStudy (anonymous):

I don't know! There are a lot of problems like these and in most of them you use combinations

Parth (parthkohli):

Combinations are used in other kinds of problems. Not the simple probability :)

OpenStudy (anonymous):

thanks for your help though :)

Parth (parthkohli):

yw :D

OpenStudy (anonymous):

alright I understand

OpenStudy (kropot72):

The hypergeometric distribution applies (sampling without replacement) \[P(2yellow)=\frac{\left(\begin{matrix}6 \\ 2\end{matrix}\right)\left(\begin{matrix}4 \\ 0\end{matrix}\right)}{\left(\begin{matrix}10 \\ 2\end{matrix}\right)}\]

OpenStudy (kropot72):

So we have: \[P(2yellow)=\frac{6\times 5\times 2}{2\times 10\times 9}\]